OpenStudy (anonymous):

Really need help understanding a couple of different problems that i will be tested over tomorrow. I'd be very gracious for the help :) 1.) simplify the cuberoot of 64x^6 2.) simplify the cuberoot of 125x^6y^9

OpenStudy (anonymous):

it'd be a lot of help to just see how you solve these 2

TheSmartOne (thesmartone):

Is this your question? \[\Large\sqrt[3]{64x^6}\] \[\Large\sqrt[3]{125x^6y^9}\]

TheSmartOne (thesmartone):

Also btw \(\bf\huge\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8 @nicoleconstance24

OpenStudy (anonymous):

yes :)

TheSmartOne (thesmartone):

So do you know that \[\Huge\sqrt[a]{b}=b^{\frac{ 1 }{ a }}\]

OpenStudy (anonymous):

yeah i vaguely remember that

TheSmartOne (thesmartone):

So lets use that formula with your first question :)

TheSmartOne (thesmartone):

\(\Large\sqrt[3]{64x^6}\) So that will become \[\Large(64x^6)^\frac{ 1 }{3 }\]

TheSmartOne (thesmartone):

Do you know that \[\Large(abc)^x= a^xb^xc^x\]

OpenStudy (anonymous):

yes

TheSmartOne (thesmartone):

Also this one :P \[\Large(abc)^\frac{1}{x}= a^\frac{1}{x} b^\frac{1}{x} c ^\frac{1}{x}\]

OpenStudy (anonymous):

not so much that one lol

OpenStudy (anonymous):

but that does make sense now that i look at it

TheSmartOne (thesmartone):

Good!

TheSmartOne (thesmartone):

And this is what it would be if there was only 2 numbers variables \[\Large(ab)^\frac{1}{x} = a^\frac{1}{x} b^\frac{1}{x}\]

TheSmartOne (thesmartone):

So lets use \[\Large(ab)^\frac{1}{x} = a^\frac{1}{x} b^\frac{1}{x}\] And figure out what \(\Large(64x^6)^\frac{ 1 }{3 }\) equals to :)

OpenStudy (anonymous):

im not sure how to simplify that though

TheSmartOne (thesmartone):

We will get to that :)

OpenStudy (anonymous):

okay good thats the part that is stumping me lol

TheSmartOne (thesmartone):

So \(\Large(64x^6)^\frac{1}{3} = {64}^\frac{1}{x} ({x^6})^\frac{1}{x}\) Have you gotten this?

TheSmartOne (thesmartone):

@nicoleconstance24

OpenStudy (anonymous):

no, i have no idea how you did that

TheSmartOne (thesmartone):

So what number. When you multiply it by itself three tims in a row equals to 64? Like 2 x 2 x2 =8. So 8 ^ 1/3 = 2

OpenStudy (anonymous):

4

TheSmartOne (thesmartone):

Correct.

TheSmartOne (thesmartone):

\[\Large(ab)^\frac{1}{x} = a^\frac{1}{x} b^\frac{1}{x}\] I had used this formula. ^

TheSmartOne (thesmartone):

Also what is